Determine the mass of 3.011 x 1023 atoms of gold.
1
Expert's answer
2020-12-03T14:05:58-0500

Now you can use proportion:

6.022 * 10^23 at. = 16 g

3.011 * 10^23 at. = x g

So x = 8g

Or you can check, how many moles is 3.011*10^23 at. - and it is 1/2, then just multiplay moles * molar mass, so you get

1/2 mol * 16 g/mol = 8 g